OpenStudy (anonymous):

the accelaration due to gravity on the moon is about one sixth of that on earth. if an object is thrown upward how much farther will it go, assuming the initial velocity is the same 5m/s?

OpenStudy (anonymous):

okay we have our formula \[v^2=u^2-2gs\] here we have v=0 u=5m/s and g=9.8/6 find s ..! :)

OpenStudy (anonymous):

Use kinematic equations: y(t) = yi + vi t + 1/2 a t^2 v(t) = vi + at vi = 5m/s yi = 0 m a = -10/6 m/s^2 for the moon and about -10 m/s^2 for earth the v(t) you will find is the velocity at the highest point the ball will travel so it will equal 0 since it is the turning point before falling back to the ground t is unknown in both cases- find it with the velocity equation using the known values for the moon case: 0 = 5 - 10/6 t t = 3 you can use t = 3 in the y position equation to find the height at that time which we know is the highest point: y(t) = 0 + 5(3) + 1/2(10/6)(3^2) y(t) = (15/2) or 7.5 m now repeat the process using the acceleration for the situation on earth and find the difference between the two heights for your answer